OpenStudy (anonymous):

Find the limit as x approaches 0:

OpenStudy (anonymous):

\[Sin^2 3x/5x^2\]

OpenStudy (anonymous):

apply l'hospital's rule a few times

OpenStudy (anonymous):

should go to 0

OpenStudy (anonymous):

I haven't learned that rule; I know I'm supposed to use the squeeze theorem

OpenStudy (sirm3d):

\[\frac{\sin^23x}{5x^2}=\frac{1}{5}\frac{\sin3x}{x}\frac{\sin 3x}{x}\]

OpenStudy (anonymous):

difficult to squeeze that one... hmmm

OpenStudy (sirm3d):

there is no need to squeeze, because it was only used to evaluate \[\lim_{x\rightarrow 0}\frac{\sin x}{x}\]

OpenStudy (anonymous):

yeah... \[\frac{ \sin ^{2}3x }{ 5x ^{2} }=\frac{ 9 }{ 5 }\frac{ \sin 3x }{ 3x }\frac{ \sin 3x }{ 3x }\]

OpenStudy (anonymous):

where is the 9 from? is the 3 just squared b/c it's sin^2?

OpenStudy (anonymous):

i know... some teachers teach that limit as the "squeeze theorem" same as the (1-cos x)/x

OpenStudy (anonymous):

sin 3x/3x needs to be that way to satisfy the limit. it's not lim (sin 3x)/x = 1 as x ->0 it's lim (sin x)/x = 1 as x -> 0. the argument of sin and the denominator have to go to 0 at the same rate.

OpenStudy (anonymous):

I understand, but where did the 9 come from?

OpenStudy (anonymous):

the 9 came from the 3x's in the denominator. if you multiply out, you'll get your original expression

OpenStudy (anonymous):

that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome!